moblern Set #1 In the So Lowin 3ง. 6 vous + ナ

1) In the following circuit, calculate i1, i2, v1, v2, and v3.
Answer #1

1) from the circuit we found that,

V1=6V, V3=-9V,

V1= V2+V3

V2=V1-V3=6-(-9)= 15V

From KCL,

i3= (6V/3 ohms)= 2Amps

i4=(-9V/3 ohms)=-3Amps

At node-a: Applying KCL

i1+5+4=2

i1=-7-mps

At node-b, Applying KCL:

4+5+i4+i2=0

9-3=-i2

i2=-6Amps

Hence, V1=6V, V2=15V, V3=-9V, i1=-7, i2=-6A, i3=2A, i4= -3A